# Issue2Test: Generating Reproducing Test Cases from Issue Reports

### v3

@@ Abstract.
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 2.1. Running Example
Title Cannot override get_FOO_display() in Django 2.2+ Description I cannot override the get_FIELD_display function on models since version $2.$ It works in version 2.1. Example: 1class FooBar(models.Model): 2 foo_bar = models.CharField(_("foo"), choices=[(1, ’foo’), (2, ’bar’)]) 3 4 def __str__(self): 5 return self.get_foo_bar_display() # This returns ’foo’ or ’bar’ in 2.2, but ’something’ in 2.1 6 7 def get_foo_bar_display(self): 8 return "something" What I expect is that I should be able to override this function.
@@ 2.3. Phase 1: Issue Comprehension
Root Cause Analysis Prompt Task: Perform root cause analysis for the following GitHub issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for paper) Propose distinct sets of root causes that could explain the issue. Integrate contextual information from the issue description. Identify mutually exclusive or independent failure scenarios.
@@
Cause 1: Handling of Choices in Inherited Models … (shortened for the paper) Cause 2: Cache or Meta-Class Processing … (shortened for the paper) Cause 3: Implementation Bug in Django’s Display Method –¿ Reason 1: There might be a bug within Django itself where the get_FOO_display() method does not account for inherited or overridden choices correctly, particularly when choices are altered at the child class level. –¿ Reason 2: The method that resolves choices within Django might be erroneously falling back to a default or initial state rather than considering the dynamic changes made by child classes, leading to incorrect value displays.
@@
Guidelines Meta Prompt Issue: [Excerpts from Django-12284, shortened for the paper] Cannot override get_FOO_display() in Django 2.2+ … Task: Provide structured guidelines for writing a self-contained and executable unit test for the issue above, which was reported in the project Django 2.2+.
@@
IssueID: django-12284 Found files: 1. django/db/models/options.py 2. django/db/models/fields/__init__.py 3. django/db/models/fields/mixins.py 4. django/db/models/base.py 5. django/db/models/query.py
@@ 2.4. Phase 2: Test Generation
Test Generation Prompt Task: Your task is to generate unit tests based on the provided GitHub issue description and source code. Each test case should be designed to fail initially, confirming the existence of a bug or unimplemented feature. GitHub Issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for brevity) Guidelines: To create an independent, executable unit test in Django, dynamically configure settings and use an in-memory database for isolation. Configure Django Settings…(shortened) Root cause analysis: The Django model’s mechanism for resolving choices in an inherited class might not correctly override … Related Source File: 1import collections.abc 2import copy 3... (shortened for paper) 4def rel_db_type(self, connection): 5 return SmallIntegerField(). 6 db_type(connection=connection) Related Test File: 1... (shortened for paper) 2class BasicFieldTests(SimpleTestCase): 3 def test_show_hidden_initial(self): 4... (shortened for paper) 5 self.assertChoicesEqual( 6 field.get_choices(include_blank=False, limit_choices_to={}),[self.bar1, self.bar2],) Expected Output: Self-contained Python test file.
@@
Error Categorization Prompt Task: Analyze the outcome of the below-failing test and classify it into one of the following: 1. Compilation Error: The test cannot be compiled. 2. Runtime Error: The test crashes during execution. 3. Assertion Failure: The test runs and fails due to an assertion. Also, tell whether the assertions or error is caused by the issue described below. GitHub Issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for paper) Test Execution Output: FAIL: test_get_field_foo_display_for_C Traceback (most recent call last): File "/testbed/tests/model_fields/test_new.py", line 43, in test_get_field_foo_display_for_C self.assertEqual(instance.get_field_foo_display(), ’output3’) AssertionError: ’C’ != ’output3’ - C + output3 Expected Response: { "error_type": "<compilation, runtime, or assertion>", "issue_error_relevance": <true or false>, "repair_steps": "<corrective actions>" }
@@ 2.5. Phase 3: Test Refinement
Test Modification Prompt(pass into fail) Task: The test case below passes even though the reported GitHub issue is unresolved. Modify the test case to ensure that it fails because of the issue below. GitHub Issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for paper) Passing Test: 1from django.conf import settings 2import django 3... (shortened for paper) 4 def test_get_field_foo_display_for_A(self): 5 instance = B(field_foo=’A’) 6 self.assertEqual(instance.get_field_foo_display(), ’output1’)
@@ 3.2.3. Examples of Successful cases
Error category: Compilation Error Explanation: The test cannot be executed because the model class ’B’ is not properly configured within an application in INSTALLED_APPS. Root cause: The model class ’B’ does not …(shortened for paper) Repair steps: 1. Add an explicit app_label to the class B in the Meta class. 2. Ensure that ’test_app’ is correctly listed in the INSTALLED_APPS configuration.
@@
Direct match: Yes Reason: The assertion failure occurs because the method get_field_foo_display() does not return the expected output for the new choice ’C’, which is directly related to the bug described in the GitHub issue… (shortened)
@@ 3.2.4. Example of Failure Cases
Direct match: Yes Reason: The assertion failure is due to the missing implementation of the feature described in the GitHub issue, which states that napoleon_use_param should also affect the ”Other parameters” section… (shortened)
@@
Direct Match: Yes Reason: The failure is expected due to the unimplemented feature in the GitHub issue, which describes an INTERNALERROR when an exception occurs in the __repr__ method. The test case is designed to check the behavior of __repr__ when an exception is raised in __getattribute__, which aligns with the issue.
Figure 15. Assertion failure reasoning for pytest-7168.
Figure 16. Venn diagram of F→P tests across approaches.
@@ 4. Discussion
5. Threats to Validity
One potential threat to internal validity is the evaluation dataset size. We evaluate Issue2Test on SWT-bench-lite, which consists of 276 issues from the larger SWT-bench dataset. While this benchmark has been widely used in prior studies, it may not fully capture the diversity of issue types encountered in practice (e.g, some projects in SWT-bench-lite have less than 10 issues). Another limitation of Issue2Test is its reliance on meta-prompting, which is particularly beneficial for popular open-source projects. However, while this approach enhances test generation accuracy, it raises concerns about Issue2Test’s ability to generalize to private repositories or lesser-known projects where LLMs lack prior exposure.
@@ 6. Related Work
Reproducible Test Case Generation. Reproducing software failures from issue reports is essential for debugging (beller:debugging:icse18), regression testing (libro:icse23), and automated program repair (rondon:agent-based-repair-google:arxiv25). In Section 3.1.4, we discussed approaches for reproducible test case generation. Additionally, BRT Agent (cheng:agentic-bug-reproduction:arxiv25) focused on generating tests from bug reports, specifically in an industrial setup. The reported success of BRT demonstrates that LLM-based test generation is beneficial not only for public issues but also for private internal issues at large companies such as Google.
@@
LLM-Based Test Generation. LLMs have been applied to test generation (feng:bug-replay-llm:icse24; libro:icse23; schafer:test-generation-using-llm:arxiv23; schafer:unit-test-generation-using-llm:tse23; cedar), with efforts to enhance quality by overcoming coverage plateaus (lemieux:codamosa:icse23), incorporating code-aware prompting (symprompt), and leveraging coverage information (coverup:arxiv24). LLMs have also been explored for augmenting existing tests (test-improvement-using-llms-meta) refines human-written test suites to improve coverage. Our work differs by focusing not on direct code generation but on synthesizing test cases specifically for reproducing issues raised on GitHub and revealing the buggy behavior related to the issue.
@@ 7. Conclusion
8. Data Availability
The paper is currently under review, and the code artifacts will be made available soon.

### v4

@@ Abstract.
Bug reproduction, test generation, large language model
@@ 2.1. Running Example
[…]
Title
Cannot override get_FOO_display() in Django 2.2+
Description
I cannot override the get_FIELD_display function on models since version 2.2. It works in version 2.1.
Example:
1class FooBar(models.Model):
2 foo_bar = models.CharField(_("foo"), choices=[(1, ’foo’), (2, ’bar’)])
3
4 def __str__(self):
5 return self.get_foo_bar_display() # This returns ’foo’ or ’bar’ in 2.2, but ’something’ in 2.1
6
7 def get_foo_bar_display(self):
8 return "something"
[…]
@@
[…]
Import Statements
1from django.conf import settings
2import django
3from django.db import models
4from django.test import TestCase
Django Settings Configuration
5if not settings.configured:
6 settings.configure( INSTALLED_APPS=[
7 ’django.contrib.contenttypes’, ’test_app’],
8 DATABASES={’default’: {’ENGINE’:
9 ’django.db.backends.sqlite3’,’NAME’:’:memory:’}})
[…]
13
14 % Helper Classes for Reproducing the Issue
[…]
17
18 % $\begin{tcolorbox}$[colframe=white, colback=white, sharp corners, boxrule=0.0pt, left=0.5pt, right=0.5pt, top=-15pt, bottom=-10pt]
19 % $\begin{whitebox}$
[…]
21class A(models.Model):
22 foo_choice = [("A", "output1"),("B", "output2")]
23 field_foo = models.CharField(max_length=254,
24 choices=foo_choice)
25 class Meta:
26 abstract = True
27
28class B(A):
29 foo_choice = [("A", "output1"), ("B", "output2")
30 ,("C", "output3")]
31 field_foo = models.CharField(max_length=254,
32 choices=foo_choice)
33 class Meta:
34 app_label = ’test_app’
Test Case and Assertion
25class ModelDisplayTests(TestCase):
26 def test_get_field_foo_display_for_C(self):
27 instance = B(field_foo=’C’)
28 self.assertEqual(instance.get_field_foo_display(),’output3’)
@@ 2.3. Phase 1: Issue Comprehension
Root Cause Analysis Prompt
Task: Perform root cause analysis for the following GitHub issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for paper) Propose distinct sets of root causes that could explain the issue. Integrate contextual information from the issue description. Identify mutually exclusive or independent failure scenarios.
@@
[…]
Cause 1: Handling of Choices in Inherited Models
… (shortened for the paper)
Cause 2: Cache or Meta-Class Processing
… (shortened for the paper)
Cause 3: Implementation Bug in Django’s Display Method
–¿ Reason 1: There might be a bug within Django itself where the get_FOO_display() method does not account for inherited or overridden choices correctly, particularly when choices are altered at the child class level.
–¿ Reason 2: The method that resolves choices within Django might be erroneously falling back to a default or initial state rather than considering the dynamic changes made by child classes, leading to incorrect value displays.
[…]
@@
Guidelines Meta Prompt
Issue: [Excerpts from Django-12284, shortened for the paper] Cannot override get_FOO_display() in Django 2.2+ … Task: Provide structured guidelines for writing a self-contained and executable unit test for the issue above, which was reported in the project Django 2.2+.
@@
[…]
IssueID: django-12284
Found files:
1. django/db/models/options.py
2. django/db/models/fields/__init__.py
3. django/db/models/fields/mixins.py
4. django/db/models/base.py
5. django/db/models/query.py
[…]
@@ 2.4. Phase 2: Test Generation
Test Generation Prompt
Task: Your task is to generate unit tests based on the provided GitHub issue description and source code. Each test case should be designed to fail initially, confirming the existence of a bug or unimplemented feature. GitHub Issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for brevity) Guidelines: To create an independent, executable unit test in Django, dynamically configure settings and use an in-memory database for isolation. Configure Django Settings…(shortened) Root cause analysis: The Django model’s mechanism for resolving choices in an inherited class might not correctly override … Related Source File: 1import collections.abc 2import copy 3... (shortened for paper) 4def rel_db_type(self, connection): 5 return SmallIntegerField(). 6 db_type(connection=connection) Related Test File: 1... (shortened for paper) 2class BasicFieldTests(SimpleTestCase): 3 def test_show_hidden_initial(self): 4... (shortened for paper) 5 self.assertChoicesEqual( 6 field.get_choices(include_blank=False, limit_choices_to={}),[self.bar1, self.bar2],) Expected Output: Self-contained Python test file.
@@
Error Categorization Prompt
Task: Analyze the outcome of the below-failing test and classify it into one of the following: 1. Compilation Error: The test cannot be compiled. 2. Runtime Error: The test crashes during execution. 3. Assertion Failure: The test runs and fails due to an assertion. Also, tell whether the assertions or error is caused by the issue described below. GitHub Issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for paper) Test Execution Output: FAIL: test_get_field_foo_display_for_C Traceback (most recent call last): File "/testbed/tests/model_fields/test_new.py", line 43, in test_get_field_foo_display_for_C self.assertEqual(instance.get_field_foo_display(), ’output3’) AssertionError: ’C’ != ’output3’ - C + output3 Expected Response: { "error_type": "<compilation, runtime, or assertion>", "issue_error_relevance": <true or false>, "repair_steps": "<corrective actions>" }
@@ 2.5. Phase 3: Test Refinement
Test Modification Prompt(pass into fail)
Task: The test case below passes even though the reported GitHub issue is unresolved. Modify the test case to ensure that it fails because of the issue below. GitHub Issue: Cannot override get_FOO_display() in Django 2.2+ Description …(shortened for paper) Passing Test: 1from django.conf import settings 2import django 3... (shortened for paper) 4 def test_get_field_foo_display_for_A(self): 5 instance = B(field_foo=’A’) 6 self.assertEqual(instance.get_field_foo_display(), ’output1’)
@@ 3.2.2. Comparison with Baselines
Figure 11. Venn diagram of F→P tests across approaches.
@@
[…]
Direct Match: Yes
Reason: The failure is expected due to the unimplemented feature in the GitHub issue, which describes an INTERNALERROR when an exception occurs in the __repr__ method.
[…]
Figure 13. Assertion failure reasoning for pytest-7168.
3.2.3. Examples of Successful cases
@@
[…]
Error category: Compilation Error
Explanation: The test cannot be executed because the model class ’B’ is not properly configured within an application in INSTALLED_APPS.
Root cause: The model class ’B’ does not …(shortened for paper)
Repair steps:
1. Add an explicit app_label to the class B in the Meta class.
2. Ensure that ’test_app’ is correctly listed in the INSTALLED_APPS configuration.
[…]
@@
[…]
Direct match: Yes
Reason: The assertion failure occurs because the method get_field_foo_display() does not return the expected output for the new choice ’C’, which is directly related to the bug described in the GitHub issue… (shortened)
[…]
@@ 3.2.4. Example of Failure Cases
[…]
Direct match: Yes
[…]
@@ 4. Discussion
Reproducibility. Our implementation and instructions for reproducing the results are available (repo).
@@ 5. Related Work
Reproducible Test Case Generation. Reproducing software failures from issue reports is essential for debugging (beller:debugging:icse18), regression testing (libro:icse23), and automated program repair (rondon:agent-based-repair-google:arxiv25; multihunk-repair:ase25). In Section 3.1.4, we discussed approaches for reproducible test case generation. Additionally, BRT Agent (cheng:agentic-bug-reproduction:arxiv25) focused on generating tests from bug reports, specifically in an industrial setup. The reported success of BRT demonstrates that LLM-based test generation is beneficial not only for public issues but also for private internal issues at large companies such as Google.
@@
LLM-Based Test Generation. LLMs have been applied to test generation (feng:bug-replay-llm:icse24; libro:icse23; schafer:test-generation-using-llm:arxiv23; schafer:unit-test-generation-using-llm:tse23; cedar), with efforts to enhance quality by overcoming coverage plateaus (lemieux:codamosa:icse23), incorporating code-aware prompting (symprompt), and leveraging coverage information (coverup:arxiv24; panta). LLMs have also been explored for augmenting existing tests (test-improvement-using-llms-meta) refines human-written test suites to improve coverage. Our work differs by focusing not on direct code generation but on synthesizing test cases specifically for reproducing issues raised on GitHub and revealing the buggy behavior related to the issue.
@@ 6. Conclusion
Acknowledgements.
This work was supported in part by the Canadian Natural Sciences and Engineering Research Council (NSERC DG), Amazon Research Awards (AWS Generative AI), the European Research Council (ERC; grant agreements 851895 and 101155832), and the German Research Foundation (DFG; projects 492507603, 516334526, and 526259073).